40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 8 18:44:01 2007, tobin, Configuration, PSL, ISS 
    Reply  Fri Nov 9 17:33:33 2007, rob, Configuration, PSL, ISS 
Message ID: 85     Entry time: Thu Nov 8 18:44:01 2007     Reply to this: 89
Author: tobin 
Type: Configuration 
Category: PSL 
Subject: ISS 
Tobin, Rob

With the Sense PD blocked, I adjusted the offset trim of the fourth stage in the ISS servo until the current shunt signal was zeroed. After this adjustment, we are able to crank the ISS gain all the way up to 30 dB without CS saturations (provided the HEPA is turned down to a very quiet level), getting about 35kHZ UGF at that gain setting. However, the current shunt mean value was still enormous.

Examining the current shunt signal on a fast scope, we saw an enormous (>2Vpp) 3.6 MHz sawtooth signal. Going up the chain of op-amps, we found that U1, as measured at the "Filter Out" testpoint, is oscillating wildly at 12 MHz (680 mVpp).
ELOG V3.1.3-